Two pipes can fill a large tank in 10 hours. One of the pipes, used alone, takes 15 hours longer than the other to fill the tank. How long would each pipe take to fill the tank alone?
Answer by stanbon(75887)

Two pipes can fill a large tank in 10 hours. One of the pipes, used alone, takes 15 hours longer than the other to fill the tank. How long would each pipe take to fill the tank alone?
-----------
Together rate = 1/10 job/hr
----
one pipe rate = 1/(x+15) job/hr
-----
other pipe rate = 1/x job/hr
===================================
Equation:
rate + rate = together rate
1/x + 1/(x+15) = 1/10
----
Multiply thru by 10x(x+15) to get:
10(x+15) + 10x = x(x+15)
20x + 150 = x^2 + 15x
x^2 - 5x - 150 = 0
(x-15)(x+10) = 0
x = 15 hr (time for the "other" pipe to fill the tank)
x+15 = 30 hr (time for the "one" pipe to fill the tank)
========================================================
Cheers,
Stan H.